A bucket has 48 liters of water in it when it is 4/7 fill. How much can it hold?
Answer by stanbon(75887)

A bucket has 48 liters of water in it when it is 4/7 fill. How much can it hold?
---
Let the bucket's capacity be "x":
Equation:
(4/7)x = 48
(1/7)x = 12
x = 84 liters
================
Cheers,
Stan H.
================